OpenStudy (anonymous):

Find the function f(x) which has the following Laplace transform: F (s) = (4s + 3) / (s^2 − 2s + 8)

OpenStudy (oblivious):

Well, the formula for inverting a laplace transform in a general case is rather cumbersome. Better to use a table of laplace transforms, which you should have been given.. A quick look online to find the equality you need: \[F(s) = (s+\lambda)/((s+\alpha)^2+\beta^2)\] Is the transform of\[\exp(-\alpha t)\(cos(\beta t)+[(\lambda-\alpha)/\beta]*sin(\beta t))\]Now you simply need to match your equation to the F(s) equation (divide num and den by 4, factor denominator leaving remainder)